# NL2KQL: From Natural Language to Kusto Query

### v2

@@ 7. Discussion And Future Work
In this work, we propose NL2KQL, an end-to-end Natural Language (NL) to Kusto Query Language (KQL) translator which capitalizes on the capabilities of LLM. This system creates a dynamic prompt composed of the schema, few-shots, selective elements of the Semantic Data Catalog, and details of the KQL syntax and its best practices to guide the model towards emitting the correct KQL. As demonstrated in ablation studies of Tables 1 and 2, the biggest boosts in performance of NL2KQL are rooted in the KQL Guide (syntax and best practices), Schema, and the Few-shots. Given the complexities of evaluating generated queries, we designed offline (based on parsing) and online (based on execution) metrics to capture the performance of the proposed method and compare it with available baselines. To the extent of our knowledge, this is the first NLQ-to-KQL solution.
@@
The benchmark evaluation dataset, which accompanies and released along this paper, has only one KQL for each NLQ. However, considering that each NLQ can be addressed with multiple equally valid KQLs, having a single ground-truth KQL can result in an under-estimation of the overall performance with respect to the offline metrics. Moreover, in the implementation of the Filter Literal Score ($Filter_{\text{lit}}$), the values were compared literatim which does not take into account variations in, for example, timestamp formats.
Among future works, we envision the trending integration of LLMs as agent-like entities for iterative interaction, refinement and error resolution to be a viable option. In an agent-like setup samples are iteratively fetched from candidate tables to extend the model’s context and encountered exceptions are used as a basis for debugging. A significant challenge associated with this agent-based iterative methodology is the increase of response times, shifting interactions away from real-time. Similarly, the Query Refiner component could have been implemented to involve instructing an LLM to regenerate the query based on the diagnostic code from the Kusto parser. While effective, this approach would have introduces additional latency due to the slow auto-regressive nature of decoder-only LLMs. Given the real-time nature of NL2KQL, response time is a concern in design decisions.
Finally, extending the NL2KQL to a chat-based text interface enables users to inquire further about the data through follow-up questions. This process generates a new KQL that incorporates the augmented context. Although this approach presents new challenges, it significantly enhances the accessibility of Kusto databases for data analysts, facilitating more intuitive interactions.
@@ B.7. JSON Parse and Join
Appendix C Query Repair Examples
Here are a few outcomes of the Query Refiner.
In the following example, the original query sought to filter data from table DeviceFileEvents based on a regex pattern applied to the ProcessCommandLine column. However, this column does not exist in this table and this query failed the parsing. Query Refiner component identified the diagnostic error, and attempted to rectify it by substituting the unavailable column with a similar column, InitiatingProcessCommandLine, present in the table DeviceFileEvent. Although this substitution enabled successful parsing of the query, the actual error for both queries is referencing the incorrect table. Consequently, both the original and repaired queries are incorrect, while the repaired query is able to be parsed successfully by Kusto parser. Due to the misidentification of the column name in the repaired query, it leads to a decrease in the Filter Column Score metric.
// Generated:
DeviceFileEvents
| where Timestamp > ago(5m)
| where FileName contains rundll32.exe
| where ProcessCommandLine matches regex ”[A-Za-z]+_666
// Fixed:
DeviceFileEvents
| where Timestamp > ago(5m)
| where FileName contains rundll32.exe
| where InitiatingProcessCommandLine matches regex ”[A-Za-z]+_666
// Ground-truth:
DeviceProcessEvents
| where Timestamp > ago(5m)
| where FileName has rundll32.exe
| where ProcessCommandLine matches regex \”[a-zA-Z]+_666\
In the following example, because of $count()$, a summarize operator is needed which is fixed by the Query Refiner.
// Generated:
DeviceEvents
| where Timestamp > ago(1d)
| count()
// Fixed:
DeviceEvents
| where Timestamp > ago(1d)
| summarize count()
// Ground-truth:
DeviceEvents
| where Timestamp > ago(1d)
| count
Appendix D Database Schema
D.1. Defender Database

### v3

@@ 7. Discussion And Future Work
In this work, we propose NL2KQL, an end-to-end Natural Language to Kusto Query Language translator which capitalizes on the capabilities of LLM. This system creates a dynamic prompt composed of the schema, few-shots, selective elements of the Semantic Data Catalog, and details of the KQL syntax and its best practices to guide the model towards emitting the correct KQL. As demonstrated in ablation studies of Tables 1 and 2, the biggest boosts in performance of NL2KQL are rooted in the KQL Guide (syntax and best practices), Schema, and the Few-shots. Given the complexities of evaluating generated queries, we designed offline (based on parsing) and online (based on execution) metrics to capture the performance of the proposed method and compare it with available baselines. To the extent of our knowledge, this is the first NLQ-to-KQL solution.
@@
The benchmark evaluation dataset, which accompanies and released along this paper, has only one KQL for each NLQ. However, considering that each NLQ can be addressed with multiple equally valid KQLs, having a single ground-truth KQL can result in an under-estimation of the overall performance with respect to the offline metrics.
Among future works, we envision the trending integration of LLMs as agent-like entities for iterative interaction, refinement and error resolution to be a viable option. Given the real-time nature of NL2KQL, response time is a concern in design decisions.
@@ Appendix A Prompt Template
# Kusto Best practices
**General Guidelines**
- Never use placeholder values
- Minimize the number of tables and columns referenced
- Only include user-provided values in your queries with the correct table and column
- All backslashes in strings must be properly escaped
- Columns can be renamed using the ‘=‘ operator
- Avoid multiple filter conditions in a single where statement
- Functions that create columns will typically be named like function_‘, function_column‘, etc.
**Operators**
- Use has_all‘, has_any‘, in~‘, and in in place of multiple and or or operators
- Consider using has or contains before parsing JSON columns to avoid expensive parsing of rows without the required keys or values
- Use case-insensitive operators when comparing string and dynamic columns
**Aggregations**
- Aggregation functions only appear in summarize statements
- When combining aggregations and scalar values, a self-join may be necessary
- \”arg\” functions do not rename columns
- Aggregation functions cannot be nested. Use two separate summarize statements instead e.g., ‘| summarize X=AggFunc1(Col1) by Col2 | summarize AggFunc2(X)‘
**Joining Tables**
- join conditions **MUST** contain only ‘==‘
- Columns that appear in multiple tables merged by a join statement have integer suffixes e.g., ColumnName, ColumnName1, etc
- Consider all kinds of join. Use innerunique to keep all columns from both tables
- Use semi-joins (‘leftsemi‘, leftantisemi‘, rightsemi‘, etc.) to only keep columns from one table
- Use anti-joins (‘leftanti‘, leftantisemi‘, rightanti‘, etc.) to exclude records that appear in a table from another
**Tips and Tricks**
- Inequality operators like between‘, ‘<=‘, etc. cannot be used to join tables. To join on a window on datetimes or other high-cardinality columns, use bin and range to define a new key
~~~kusto
Table1
| extend Key=bin(TimeColumn, WindowSize)
| join (
Table2
| mv-expand Key=range(bin(TimeColumn - WindowSize, WindowSize / 2), bin(TimeColumn, WindowSize / 2), WindowSize / 2) to typeof(datetime)
) on Key
| where TimeColumn1 - TimeColumn between (0m .. WindowSize)
- To compare an aggregated and unaggregated version of the same column, you will need to use a self-join
~~~kusto
Table
| summarize VarName=AggFunc(Col1) by Col2
| join Table on \$left.Col2 == \$right.Col2
| where Col1 < VarName
- Some columns may be blank or dependent on other column values. Use a self-merge to utilize more than one of these columns at a time
~~~kusto
Table
| where IndependentColumn == Value1
| where isnotempty(DependentColumn1)
| join (
Table
| where IndependentColumn == Value2
| where isnotempty(DependentColumn2)
) on JoinColumn
| project JoinColumn, IndependentColumn, DependentColumn1, DependentColumn2
@@ Appendix B KQL Forms
<TableExpression>
| <operator> [parameters]
| <operator> [parameters]
// Data Retrieval
TableName
| take 100
// Filtering
TableName
| where ColName == Value
// Aggregation
TableName
| summarize Count = count() by ColName
// Sorting
TableName
| order by ColName asc
@@ B.7. JSON Parse and Join
Appendix C Database Schema
C.1. Defender Database